Please check my work:
Form each of the following:
A linear equation in one variable: X + 3 = -5
A linear equation in two variables: y = 5x - 3
A quadratic equation: 3x2 + x - 2 = 0
A polynomial of three terms: 7x2y3 + 4x − 9
An exponential function: 32 = 9
A logarithmic function: log39 = 2

Answer by stanbon(75887)

Please check my work:
Form each of the following:
A linear equation in one variable: Y + 3 = -5
A linear equation in two variables: y = 5x - 3
A quadratic equation: 3x2 + x - 2 = Y
A polynomial of three terms: 7x^2 + 4x − 9 = y
An exponential function: y = 3^x +2
A logarithmic function: y = 3 + log(x-2)
=============================================
Cheers,
Stan H.